### v2

# Finer: Investigating and Enhancing Fine-Grained Visual Concept Recognition in Large Vision Language Models

Abstract
Recent advances in instruction-tuned Large Vision-Language Models (LVLMs) have imbued the models with the ability to generate high-level, image-grounded explanations with ease. While such capability is largely attributed to the rich world knowledge contained within the Large Language Models (LLMs), our work reveals their shortcomings in fine-grained visual categorization (FGVC) across six different benchmark settings. Most recent state-of-the-art LVLMs like LLaVa-1.5, InstructBLIP and GPT-4V not only severely deteriorate in terms of classification performance, e.g., average drop of 65.58 in EM for Stanford Dogs for LLaVA-1.5, but also struggle to generate an accurate explanation with detailed attributes based on the concept that appears within an input image despite their capability to generate holistic image-level descriptions. In-depth analyses show that instruction-tuned LVLMs exhibit modality gap, showing discrepancy when given textual and visual inputs that correspond to the same concept, preventing the image modality from leveraging the rich parametric knowledge within the LLMs. In an effort to further the community’s endeavor in this direction, we propose a multiple granularity attribute-centric evaluation benchmark, Finer, which aims to establish a ground to evaluate LVLMs’ fine-grained visual comprehension ability and provide significantly improved explainability. [^1] We release the dataset and code at: https://github.com/wjdghks950/Finer
1 Introduction
Figure 1: Current state-of-the-art LVLMs exhibit strong zero-shot downstream task solving abilities (e.g., image captioning, VQA, reasoning). However, when prompted to classify the fine-grained concepts, most of them fail to distinguish them into finer categories. Fine-grained classification prompt here is omitted for brevity.
The surge of interest and development of Large Vision-Language Models (LVLMs) that are able to generate image-grounded texts have seen significant progress. Models such as InstructBLIP (Dai et al., 2023) and LLaVA (Liu et al., 2023b, a) have consistently exhibited strong zero-shot capability in generating image captions, visual reasoning and textual descriptions, and even leveraging external knowledge for complex question answering tasks Marino et al. (2019); Schwenk et al. (2022). Such results across diverse benchmarks indicate that these models, most of them being built on large language models (LLMs) like Vicuna (Chiang et al., 2023), Flan-T5 (Chung et al., 2022), Llama (Touvron et al., 2023), are already equipped with the ability to simultaneously leverage the interplay between the textual knowledge acquired during pre-training and the image understanding ability acquired during instruction-tuning. Furthermore, all of these models exhibit strong zero-shot task transferability to multiple downstream tasks.
Conventionally, in the image classification field, many previous works on fine-grained visual classification (FGVC) (Wei et al., 2022b; Diao et al., 2022; Zhu et al., 2022; Yang et al., 2022; Wang et al., 2022) sought to accurately classify diverse images ranging from different types of birds, plants and animals Van Horn et al. (2015, 2018) and artificial objects such as cars Krause et al. (2013) and aircrafts Maji et al. (2013). In this work, we investigate whether state-of-the-art LVLMs can combine their image understanding ability and rich knowledge acquired from the Web to handle FGVC tasks. To our surprise, while the LVLMs perform almost perfectly, e.g., 98.43 for LLaVA-1.5 (13B) on iNaturalist, at the superordinate-level granularity (e.g., birds, jets), their classification abilities do not extend to the coarse and finer-grained concepts (e.g., bald eagle, F-22 Raptor), exhibiting substantially deteriorated classification performance (§3.2); 46.91 for coarse-level and 1.56 for fine-level categories on iNaturalist. Our empirical analyses of these models reveal that these seemingly strong models suffer from modality gap (Liang et al., 2022). We empirically demonstrate that such discrepancy stems from LVLMs projecting visual and textual modalities disparately, limiting them from leveraging rich parametric knowledge, obtained during the textual pre-training stage of LLMs, to infer fine-grained concepts. Our results also show that such constraints lead to diminished fine-grained comprehension of the image, preventing these models from generating high-quality, finer granularity description of the concepts that appear within an image.
We also present an attribute-centric and multiple granularity image classification benchmark, Finer. Our benchmark broadens the scope of six widely-used FGVC benchmarks like iNaturalist Van Horn et al. (2018) and FGVC-Aircrafts Maji et al. (2013) by (i) generating multiple granularity levels for fine-grained visual concept recognition, and (ii) constructing a visual, concept-indicative attribute sets per fine-grained concept to measure the ability of LVLMs to generate fine-grained concept descriptions given an image. To summarize, our contributions include:
  • We highlight the lack of fine-grained image comprehension ability of instruction-tuned LVLMs across various real-life objects. To the best of our knowledge, we are the first to explore FGVC as an evaluation criteria for these models.
  • We highlight the persistence of modality gap within the LVLMs by conducting an extensive per-modality-based probing, revealing the discrepancy in how the two modalities are processed by these models.
  • We construct a novel attribute-centric, multi-granularity benchmark on FGVC datasets to open up a new direction for future works to measure LVLMs’ modality gap and their fine-grained image understanding capability. Fine-tuning on Finer proves beneficial by closing the modality gap and improving FGVC accuracy for held-out datasets.
2 Related Work
2.1 Instruction-tuning of Large Vision- Language Models
The recent advances in LVLMs such as LLaVA (Liu et al., 2023a, b), BLIP-2 Li et al. (2023), InstructBLIP (Dai et al., 2023), and closed-source models like GPT-4V OpenAI (2023); Yang et al. (2023) have brought to our attention their zero-shot task solving abilities, especially in downstream benchmarks such as Visual Question Answering (VQA), reasoning and image captioning, all of which require output generation conditioned on the extensive knowledge of the real-world. Their reasoning and captioning abilities, based on an intricate interplay between their parametric knowledge and observation of images, result in sensible outputs. However, none of them evaluates their visual comprehension abilities to identify fine-grained objects that appear within an input image.
2.2 Vision-Language Alignment Models
Alignment-based VLMs like CLIP Radford et al. (2021) and GLIP (Li et al., 2022) rely on minimizing the contrastive objective between the visual and textual representations. A recent analytical work (Liang et al., 2022) on CLIP-like models reveals that there is a modality gap between the text and visual modalities, which imposes substantial implications on downstream task performances. Our work shows that such modality gap persists even after the image representations have been projected onto the textual representational space, which we hypothesize in this work as the underlying reason behind the substantial degradation in FGVC.
Figure 2: State-of-the-art instruction-tuned LVLM zero-shot performance on fine-grained classification. All the models exhibit strong classification capabilities when prompted to classify superordinate-level (e.g., birds, cars) and coarse-grained categories(e.g., owls, SUVs), but exhibit significant deterioration in performance when prompted to categorize more fine-grained categories on the same images. The gold tags for coarse- and fine-grained classifications denote the use of gold labels from the parent category in the prompt.
2.3 Fine-grained Visual Categorization
Previous works approach FGVC with masked image modeling (He et al., 2022; Ryali et al., 2023) or incorporating concept-related meta-information (e.g., location, spatio-temporal prior, text description) (Diao et al., 2022) or LLM-generated textual descriptions of concepts (Menon and Vondrick, 2023). However, learning of fine-grained visual categories in LVLMs and their ability to elaborate on the fine-grained details of the input image via text generation is yet to be explored. Furthermore, recent works in FGVC have also shown to disregard fine-grained details of images (Krojer et al., 2022) and videos Wang et al. (2023), work poorly on downstream tasks involving localization (Ranasinghe et al., 2023; Zhong et al., 2022), object attributes (Yuksekgonul et al., 2022) and actions Wang et al. (2023).
Figure 3: Fine-grained classification pipeline. At each granularity level, a predicted output from LVLM is injected into the next level prompt. (1) Superordinate-level prompt is used to predict the highest-level category (e.g., bird). (2) Coarse-level prompt is subsequently fed with the predicted output and fed back to the LVLM to generate the next output (e.g., parrot), and (3a) and (4) follow the same steps. (3b) illustrates AttrSeek, an alternative path for (3a), wherein the model generates the attributes first.
3 Fine-Grained Image Understanding in Vision-Language Models
We first evaluate the fine-grained visual categorization performance of five different instruction-tuned baselines on six different FGVC benchmarks. This section elaborates on the details of the experimental setup and models investigated in this work.
3.1 Evaluation Settings
Datasets
Covering a wide range of real-world objects over various categories, existing FGVC benchmarks provide richly annotated set of image-concept pairs. As shown in Figure 2, we use iNaturalist-2021 (Van Horn et al., 2018), FGVC-Aircrafts (Maji et al., 2013), Stanford Dogs (Khosla et al., 2011), Stanford Cars (Krause et al., 2013), NABirds (Van Horn et al., 2015), and CUB-200-2011 (Wah et al., 2023). For each dataset, we divide the ground-truth concept label into three levels of granularity: superordinate, coarse and fine, as defined in a previous work (Hajibayova, 2013). Superordinate level refers to the highest taxonomic concepts (e.g., bird, car), coarse level refers to the lower-level granularity concepts (e.g., parrot, SUV), and fine level refers to the lowest, finer-level granularity (e.g., owl parrot (Strigops habroptila), Hyundai Santa Fe 2018). We discuss each benchmark and the construction of superordinate and coarse-grained labels in detail in Section 5.
Metrics
We assess the accuracy of the generated concept labels given an image and granularity-specific prompt asking the model to figure out the correct category the concept in the image belongs to. Following previous works on concept classification using auto-regressive models, we employ F1 and Exact Match (EM) scores; note that the EM score used in this work is a modified EM score that parses a sentence and considers the output label correct if the ground-truth label string exists within a pre-defined maximum number of tokens, $m$ (in this work, we set this value to $m=20$).
Models
The models used in this work are as follows: LLaVA-1.5 (Liu et al., 2023b, a), InstructBLIP (Dai et al., 2023), BLIP-2 (Li et al., 2023), and GPT-4V; the hyperparameter setting for each model is in Appendix A.1. The open-sourced models like LLaVA-1.5 and InstructBLIP follow a generic pipeline of transforming an input image $\mathbf{X_{v}}$ with a frozen vision encoder such as CLIP ViT-L/14 Radford et al. (2021) into an encoded image representation $\mathbf{Z_{v}}$. Then, these models either project $\mathbf{Z_{v}}$ into the language representation space through a learned projection layer $\mathbf{W}$, which becomes $\mathbf{H_{v}}=\mathbf{W}\cdot\mathbf{Z_{v}}$ as in Liu et al. (2023b), or attend over $\mathbf{Z_{v}}$ with learnable queries $\mathbf{Q}$ as in Li et al. (2023); Dai et al. (2023). Such transformed visual representations interact with $\mathbf{X_{instruct}}$, a language instruction, which attends over the image representations (or queries) within the self-attention layers of the LLM component, with the model generating the final output sequence.
3.2 Brittleness of Vision-Language Models
Zero-shot Model Performance on FGVC
To evaluate the fine-grained image categorization ability of LVLMs, we measure their classification performance per granularity by prompting the models to generate the correct label for a given concept image as shown in Figure 2 . As elaborated in Figure 3, we assess the models’ classification ability across three different granularity levels. In Figure 2, we evidence significant deterioration in terms of classification performance across all the five baselines, with some even reaching near 0% in EM score. While models do perform very well for superordinate-level categories, with some reaching 100% in EM, the finer granularity leads to worse classification performance. In terms of model size, larger models like LLaVA-1.5 (13B) and GPT-4V tend to perform better than smaller models like the 7B variants. For InstructBLIP, the 7B version performs better than the 13B version by a large margin, with the 13B version exhibiting less capable instruction-following ability than the 7B one, potentially due to 7B variant being less prone to overfitting and exhibiting efficiency in simple tasks like classification.
Models Superordinate Coarse Fine
LLaVA-1.5 (13B) 96.872 40.625 1.562
 + CoT (0-shot) - 35.910 4.687
 + CoT (3-shot) - 34.925 2.159
 + AttrSeek - 35.937 1.562
InstructBLIP (13B) 98.437 50.221 3.714
 + CoT (0-shot) - 31.805 1.238
 + CoT (3-shot) - 30.591 0.919
 + AttrSeek - 30.177 3.020
GPT-4V 100.00 83.115 18.752
 + CoT (0-shot) - 93.750 20.312
 + CoT (3-shot) - 95.283 24.389
 + AttrSeek - 95.331 53.125
Table 1: Elicitive prompting results (%) on iNaturalist. Prompting techniques like Chain-of-Thought (CoT) on the fine-grained classification (Fine) is unable to improve performance in open-source models.
Elicitive Prompting for FGVC
LLMs like Vicuna (Chiang et al., 2023) and LLama (Touvron et al., 2023) used as textual reasoning components of VLMs are known to perform better when presented with elicitive prompts like Chain-of-Thought (CoT) (Wei et al., 2022a) that improve model’s reasoning ability. A unifying thought along this line of prompting techniques is to break down a complex problem into a sequence of sub-problems, i.e., divide-and-conquer. Inspired by these prompting techniques, we propose and evaluate our prompting technique for FGVC, AttrSeek. In this simple prompting strategy, we first (i) prompt the VLMs to generate the most distinctive physical attributes visible in the concepts in an image, and (ii) feed the generated set of attributes along with the concept-asking prompt for final prediction as shown in Figure 3. However, in Table 1, we do not see much improvement in terms of fine-grained concept classification in the open-source models, with neither CoT nor AttrSeek enhancing the classification performance. However, there is a substantial increase in fine-grained performance for GPT-4V when prompted with first with the attribute-seeking prompt and then predicting the end concept. This result suggests that LLaVA-1.5 and InstructBLIP, while they exhibit strong image captioning and reasoning ability, is limited in terms of image grounded attribute understanding even when provided with elicitive prompts. This result also suggests that open-source VLMs may lag behind in instruction-following abilities in comparison to GPT-4V For additional details on few-shot prompting, see Appendix A.3.
4 Modality Gap: Discrepancy Between Textual and Visual Modalities
We first hypothesize that the lack of zero-shot concept classification ability of LVLMs is mainly due to the modality gap of textual and visual inputs, preventing the models from leveraging the existing concept-related parametric knowledge when an image of a concept is given. In this section, we aim to delve into how these models process visual and textual modalities by exploring how well they perform when given only textual descriptions of a concept (§4.1) and how they accurately elaborate what they see in a given image (§4.2).
Base Model LLM Input Modality Similarity Measurement Metrics ROUGE-1 ROUGE-2 ROUGE-L BertScore AlignScore LLaVA-1.5 (Liu et al., 2023a) Vicuna-7B Text 22.343 11.023 21.161 85.429 18.935 Image 17.468 10.049 17.314 84.499 15.509 Vicuna-13B Text 20.616 10.902 19.647 85.944 15.746 Image 19.661 11.334 19.114 85.704 13.319 $\ \ \Delta$ Avg. 2.920 0.703 2.189 0.585 2.926 InstructBLIP (Dai et al., 2023) Vicuna-7B Text 18.987 9.192 16.577 83.102 15.002 Image 16.018 9.788 11.258 82.681 10.186 Vicuna-13B Text 13.731 7.541 16.446 80.549 13.884 Image 10.949 6.009 9.583 79.277 10.191 $\ \ \Delta$ Avg. 2.875 1.064 6.091 0.846 4.254 BLIP-2 (Li et al., 2023) Flan-T5-XL Text 20.722 10.511 20.118 85.859 15.827 Image 14.107 6.173 13.376 81.343 11.673 Flan-T5-XXL Text 23.174 12.482 21.891 86.720 19.680 Image 18.398 9.490 16.237 84.190 15.294 $\ \ \Delta$ Avg. 5.695 3.664 6.197 3.522 4.270 GPT-4V (OpenAI, 2023) Text 24.320 10.179 22.748 87.457 10.524 Image 22.675 8.812 20.477 85.495 5.067 $\ \ \Delta$ Avg. 1.644 1.367 2.271 1.961 5.456
Table 2: Measuring the modality gap via textual similarity against the Web-extracted concept attributes against the LVLM-generated attributes for the iNaturalist subset in Finer. The discrepancy between the attributes generated from Text-only input and Image-only input indicate that VLMs treat the two modalities of the same concept differently. $\ \ \Delta$ Avg. indicates the average difference between the Text and Image outputs against the reference.
4.1 Probing for Concept-Related Parametric Knowledge
Figure 4: Model Performance on Text-only vs. Image-only Inputs. LLaVA-1.5 (7B and 13B), when provided only the textual information (7B-, 13B-Text) related to the ground-truth concept, outperforms the image-only input (7B-, 13B-Image) counterpart.
With the drastic performance degradation discussed in Section 3.2, we first need to verify whether concept-related knowledge already exists within the parametric knowledge space of the LVLMs to make sure that the models have already acquired concept-related knowledge necessary for zero-shot classification during the LLM pre-training stage. While we may assume that such knowledge has already been captured during pre-training of the LLM component on the Web-crawled concept documents, it is crucial to elucidate whether the knowledge exists within its parameters and whether the image representation is able to accurately capture such attributes of a concept that enables the models to infer the correct end concept. In this experiment, we measure the classification performance of the models with two different input types: (i) text-only input that consists of a concept’s external physical, descriptive attributes (details about the attribute extraction in §5), and (ii) image-only input that consists of the image of the concept; the generated outputs are measured in the same way as in Section 3.1. The text-only input, $\mathbf{X_{txt}}=[I;Attr;C]$, is composed of Instruction ($I$), web-extracted attributes ($Attr$), and coarse-grained labels ($C$); the output is the fine-grained concept label. Note, for fair comparison, we also include $C$ as input for image-only probing. In Figure 4, we show the results across three different datasets, iNaturalist, FGVC-Aircraft and Stanford Dogs. The results show that even with text-only input that contains the detailed physical attributes of a concept, LVLMs are capable of solving fine-grained visual classification, outperforming the image-only input, which is the model’s zero-shot performance on fine-grained classification. Additionally, we can also see that the larger the model size, the better the performance; this could be related to the amount of parametric knowledge that resides in a model’s LLM module.
4.2 Measuring the Modality Gap with Attribute Generation
Having observed the discrepancy between the visual and textual modalities, we now analyze whether LVLMs can observe and tell visually grounded physical attributes of an input image. We construct a set of Web-extracted concept attributes from Wikipedia documents (further elaborated in detail in Figure 5; §5) to be used as the reference texts. Then, we prompt the VLMs to generate a set of external, discriminative physical attributes of a concept when given either an image or text input (see the detailed prompts in Appendix 10); the text input refers to the concept label of the text along with a prompt that asks for the concept’s external physical descriptions. The textual similarities between the LVLM-generated attributes and the Web-extracted attributes are compared using five different scoring metrics that span both the token-level overlap and NLI model-based semantic similarity measure: ROUGE-1, 2, L (i.e., Longest-substring match), BertScore (Zhang et al., 2019), and AlignScore (Zha et al., 2023). Both the model-generated attributes and web-extracted attributes are linearized for textual similarity comparison; we detail the linearization process in Appendix A.2.
As shown in Table 2, the text-only inputs show greater textual similarity to the Web-extracted reference attributes, indicating that while the attribute-specific knowledge acquired is being used by the textual modality, the visual modality does not leverage such knowledge to a degree that matches the textual modality. Aside from the modality discrepancy, a plausible interpretation is that these models fall short on accurately focusing on specific parts of the image, diluting away the fine-grained details such as stripes or patterns (Table 4) for a higher-level, semantic understanding of the image.
5 Finer: Fine-Grained Concept Categorization Benchmark
To facilitate the research for fine-grained image understanding in LVLMs, we propose a multiple granularity benchmark, Finer. Our benchmark intends to evaluate and improve the model capabilities in terms of improving the interplay between image understanding and textual latent knowledge of the concepts that appear within the images, based on the observation from previous sections (§3.2, 4)
Figure 5: Depiction of the Finer benchmark construction pipeline. Following the aggregation of the six benchmarks in the FGVC domain, concept attributes and concept images are retrieved from Wikipedia.
Dataset Name Total # of Instances (Test / Training) # of Superordinate Categories # of Coarse Categories # of Fine Categories # of Images Annotation Granularity Hierarchy Partial Attribute Annotation iNaturalist-2021 100K / 2.6M 11 1,103 10K 3,286,843 × CUB-200-2011 5,794 / 5,994 1 59 200 11,788 × FGVC-Aircraft 3,333 / 6,667 1 30 100 10,000 × Stanford Dogs 8,580 / 12,000 1 - 120 20,580 × × NABirds 24,633 / 2,929 1 146 404 48,562 Stanford Cars 8,144 / 8,041 1 7 196 16,185 × × FINER 372K/2.63M 16 1,416 11,171 3,393,958
Table 3: Overview of the FGVC benchmarks. Our Finer dataset demonstrates richer set of attributes per concept that enables the evaluation of fine-grained image comprehension. We also augment the benchmarks without Granularity Hierarchy with Superordinate Categories and Coarse Categories.
5.1 Dataset Construction
We construct the benchmark based on six different FGVC datasets from the computer vision domain: iNaturalist-2021 (Van Horn et al., 2018), FGVC-Aircrafts (Maji et al., 2013), Stanford Dogs (Khosla et al., 2011), Stanford Cars (Krause et al., 2013), NABirds (Van Horn et al., 2015), and CUB-200-2011 (Wah et al., 2023). These datasets span a wide range of real-life objects such as airplanes, dogs, plants, birds, mammals and cars, challenging the models to cover a variety of fine-grained concepts.
We first crawl all Wikipedia documents and their main images via a search API  [^2] https://github.com/goldsmith/Wikipedia in Python. The collected documents are then fed into GPT-4V (OpenAI, 2023) with an instruction to extract the document-grounded external, visible attributes of a concept. To briefly elaborate, we divide the extracted attributes into two different types: (i) required and (ii) likely attributes. The "required" attributes are the external, concept-indicative attributes that can be used for concept identification, e.g., blue-tailed hawks with black thorax with a broad apple green stripe, while the "likely" attributes are attributes that may co-occur with the concept, but may not be helpful in distinguishing the end concept, e.g., blue-tailed hawks inhabit trees; we provide the likely attributes as an additional meta-information for each concept. We also populate the dataset with superordinate and coarse-level concept labels. For example, as shown in Figure 5, we assign "Airplane" as the superordinate-level label and "Lockheed Martin" as the coarse-level label since FGVC-Aircrafts dataset provides a concept granularity hierarchy, e.g., Boeing $\rightarrow$ Boeing 707 $\rightarrow$ Boeing 707 MAX. However, datasets like Stanford Dogs do not provide neither the superordinate nor coarse-level labels for their fine-grained concepts. We therefore few-shot prompt GPT-4V to generate the coarse-grained labels and manually inspect their validity. We provide the dataset statistics in Table 3 and prompts for dataset generation in Appendix B.3, B.4.
Image (Concept Name) Attributes (Text-only) Attributes (Image-only) Web-Extracted Attributes Dragonfly | Orthetrum Triangulare [Uncaptioned image] Blue tail ; Broad wings; White throat;Dark head; Banded tail Elongated body; Two pairs of wings; Six legs, Compound eyes , Segmented abdomen Dark face; Bluish eyes; Black thorax with a broad apple green stripe on both sides; Black segments 1-2 and 8-10 on the abdomen; Remaining segments of the abdomen pruinosed with azure blue Pinscher | Affenpinscher [Uncaptioned image] Black, gray, silver, or tan fur ; Rough, shaggy coat; Distinct "monkey-like" facial expression; Prominent chin and jaw; Small, round, dark eyes; Ears set high and usually cropped to a point Presence of fur; Four-legged stance; Distinct muzzle with a nose; Visible tail ; Ears that are either erect or floppy Harsh rough coat when not clipped; Shaggier coat over the head and shoulders forming a mane; Shorter coat over the back and hind quarters; Notable monkey-like expression; Coat is harsh and wiry in texture when properly maintained Embraer | Embraer ERJ 145 [Uncaptioned image] T-tail design ; Straight wing design with no winglets; Mounted engines on the rear fuselage; Small cockpit windows compared to the body size; Three sets of landing gear Fixed wings on either side of the fuselage ; Cockpit windows at the front of the fuselage; Jet engines, either under the wings or mounted on the rear of the fuselage; Landing gear with wheels for taking off T-tail configuration; Two rear-mounted Rolls-Royce AE 3007 series turbofan engines; Straight wing with no winglets; Narrow, tube-like fuselage; Short, nearly oval-shaped passenger windows.
Table 4: Qualitative Analysis of the Text-only generated attributes and Image-only generated attributes from GPT-4V. The generated attributes based on the Text-only and Image-only dataset exhibits notable discrepancy. Juxtaposed with the Web-Extracted Attributes, Image-only attributes are less specific and explain more generic physical descriptions rather than those discriminative of the concept, while Text-only attributes capture finer-grained details of the concept. We provide the coarse-level and fine-level labels along with the images, and the bounding boxes are only drawn to match the highlighted text and are not part of the dataset.
5.2 Qualitative Analysis
Models CUB Stanford Dogs NABirds
LLaVA-1.5 2.744 30.990 2.078
LLaVA-1.5 (iNat) 2.709 33.473 1.562
LLaVA-1.5 (iNat-Attr) 5.177 32.354 5.728
Table 5: LLaVA-1.5 (7B) exhibits enhanced fine-grained performance after learning to generate textual attributes based on image input. (iNat) is fine-tuned directly on the concept label and (iNat-Attr) is fine-tuned to generate the attributes first and then predict the concept label.
In Table 4, we provide the Web-extracted attributes from our dataset, along with model-generated attributes as a case study on the modality gap phenomenon from Section 4. The attributes are from GPT-4V, but note that the generated attributes from other models such as LLaVA-1.5, all exhibit a similar trend. When provided with image-only input, the model generates a set of attributes that pertain to the image, e.g., for dragonfly it describes its elongated body and two pairs of wings. However, the attributes from image-only input are more generic compared to those from text-only input. In other words, these attributes are not discriminative enough to identify the input image as a specific concept category. This again suggests that these models not only fail to properly observe the fine-grained details of a concept, but fail to leverage the knowledge contained within its own parameters as can be seen from the outputs of text-only inputs.
5.3 Zero-Shot Transferability of Learning to Generate Attributes
To substantiate the effectiveness of the Web-extracted attributes in Finer, we construct an instruction-tuning mixture with the AttrSeek pipeline using the iNaturalist subset on top of LLaVA-1.5’s instruction-tuning dataset (see Appendix A.4). We fine-tune LLaVA-1.5 (7B) and evaluate the fine-tuned model on three different held-out datasets. To briefly elaborate, each instance in our mixture follows three steps: (i) Ask the model to identify the course-grained concept category, (ii) generate the visual attributes of the concept, (iii) predict the fine-grained concept label. In Table 5, we see that the model outperforms simple fine-tuning and zero-shot classification performance on the three held-out datasets. This result suggests that in FGVC, instruction-tuning LVLMs by first forcing the model to look at the apparent external, visible attributes in the image by explicitly generating the textual attributes, and then subsequently performing classification improves the model’s performance. For additional details on the instruction-tuning on our mixture, refer to Appendix A.
6 Discussion and Conclusion
In this paper, we investigate the substantial lack of fine-grained visual comprehension (FGVC) ability of instruction-tuned LVLMs. Our empirical results reveal that significant modality gap exists between the visual and textual modalities, preventing the models from effectively leveraging textual knowledge for zero-shot fine-grained visual classification. Furthermore, our results also demonstrate that while LVLMs are able to describe generic attributes of a concept, they exhibit hallucination and are unable to ground their explanation on the concept present within an input image. To establish zero-shot FGVC research among LVLMs and facilitate their in-depth understanding of apparent attributes, we proposed a novel multiple granularity, attribute-centric benchmark, Finer.
7 Limitations
Intra-Concept Variance in Images
Images of a single concept can appear in various different forms. Some images may have the whole view of the concept, while other images may have certain parts of the concept (e.g., legs, wings) partially occluded. The attributes collected per concept in Finer are constructed to be visually-grounded based on the textual attributes extracted from Web documents that pertain to these concepts; nonetheless, such intra-concept variance among images may render an attribute obsolete for certain images. Other problems such as low-quality images may also lead to this issue. In future work, exploring the visual "ground-ability" of each attribute through image-text retrieval may be a plausible approach to identifying both the most discriminative attributes that pertains to an image and the fine-grained concept label.
Selection of Baseline Models
While our work covers LVLMs that receive image and text as input, there are other VLMs such as Kosmos-2 (Peng et al., 2023), Shikra (Chen et al., 2023) and Ferret (You et al., 2023) out there, that receive bounding-box annotated images as input. However, our work only deals with un-marked image and prompt inputs, since the objective of this research is to see whether LVLMs without any referring markings (e.g., bounding boxes), can ground their generative capabilities on the input image.
Acknowledgement
We thank the anonymous reviewers for their suggestions and comments. We also would like to thank Ansel Blume, Derek Hoiem and Carl Vondrick for their ideas, feedback and support for this work. This research is based upon work supported by U.S. DARPA ECOLE Program No. #HR00112390060. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix A Experiment Details
In this section, we elaborate in detail the experimental settings of our work, including the hyperparameter settings of the large vision-language models (LVLMs), and the large language models (LLMs), which are used as a major driving block of these LVLMs. In addition to the experiment settings, we also provide details on the training mixture used in 5.3 and the fine-tuning settings of LLaVA-1.5 (7B).
A.1 Hyperparameter Settings of Vision-Language Models
Inference Settings
Our work deals with the inference-time fine-grained image understanding abilities of VLMs. The hyperparameter settings provided in the table are only for inference time (§3.2, 4, 4.1) and are not to be used for fine-tuning.
Hyperparameter LLaVA InstBLIP BLIP-2 GPT-4
max_seq_len 256 256 256 256
top_p 1.0 0.95 0.95 1.0
temperature 0.2 0.75 0.75 0.2
Table 6: Hyperparameters of the Vision-Language Models (VLMs) studied in this work. LLaVA refers to LLaVA-1.5 and InstBLIP refers to the InstructBLIP model. All the VLMs use the same hyperparameter settings for both the 7B and 13B variants; the same applies to Flan-T5-XL and Flan-T5-XXL models.
For LLaVA-1.5 models, we use Vicuna-v1.5, and for InstructBLIP models, we use Vicuna-v1.1; these are the original settings of both of these models.
Fine-Tuning Settings
In Table 7, we also provide the hyperparameter settings of the instruction-tuned LLaVA-1.5 (7B) in Section 5.3. The fine-tuning was conducted with LoRA (Hu et al., 2021) on 4 V100 (16G) for approximately 28 hours on the training mixture, which is elaborated in detail in Section A.4.
Hyperparameter LLaVA-1.5 (7B)
max_seq_len 256
top_p 1.0
temperature 0.2
lora_r 128
lora_alpha 256
gradient_accum. 16
batch_size 1
learning_rate 2e-4
lr_schedule cosine decay
optimizer AdamW
weight_decay 0.0
warmup_ratio 0.03
bf16 False
fp16 True
Table 7: Hyperparameters of LLaVA-1.5 (7B) for Instruction-tuning on the Finer training mixture.
A.2 Additional Details on Modality Gap Experiment
Using LLMs for Text-Only Setting
In Section 4, we experiment on the LVLMs’ ability to generate textual attributes based on either the text-only input or the image-only input. The caveat of using InstructBLIP and BLIP-2 models is that this models do not simply allow text-only input like LLaVA-1.5 or GPT-4V. We therefore opt to use their LLM components, Vicuna-7B and -13B, Flan-T5-XL and -XXL to generate the attributes for the text-only input. The validity of this setting holds since our goal is to compare the modality gap via how much the model stores the concept-related knowledge within its parameters.
Metrics
For ROUGE, we used the Python’s ROUGE API [^3] https://pypi.org/project/rouge/ for calculation, and we only use the F1 score in this work. As for the model-based metrics, for BertScore (Zhang et al., 2019) we use the bert-base-uncased and for AlignScore (Zha et al., 2023) we use roberta-large. These model-based metrics are state-of-the-art models for textual faithfulness evaluation, making them fit to evaluate both the faithfulness and textual consistency of the generated text.
Linearization of the Web-Extracted Attributes
In this section, we elaborate on the linearization process of the Web-extracted attributes, which are used as ground-truth reference texts in Section 4. The linearization of the attributes is a simple process of concatenating the fine-grained concept name and the generated attributes of the concept. For example, for a dragonfly named Orthetrum Triangulare, we construct a linearized string that says, Orthetrum Triangulare exhibits blue eyes, black thorax with a broad apple green stripe. The format is as follows: [<concept-name>; exhibits; attribute-1; attribute-2; ...; attribute-j]; the list is converted into a single sentence for evaluation;. The same process applies to both the LVLM-generated attributes and Web-extracted attributes to compare their textual similarity in Section 4.
A.3 Enabling Multiple Image Inputs in LVLMs for Few-Shot Prompting
The three open-sourced LVLMs investigated in our work, LLaVA-1.5 (Liu et al., 2023b, a), InstructBLIP (Dai et al., 2023) and BLIP-2 (Li et al., 2023) were all pre-trained and instruction-tuned based on a single image input and a consecutive sequence of pertinent textual instructions. Nonetheless, these models are capable of receiving multiple images given its input format. For LLaVA-1.5, we simply provide the few-shot ($k$) examples sampled from iNaturalist dataset in by interleaving $k$ <image> tokens along with the input prompt and their ground-truth concept labels. For InstructBLIP and BLIP-2, since they require attending over the input image via cross-attention with a pre-defined set of query embeddings, we first embed each of the $k$ few-shot sample images with the vision encoder. Then, the image embeddings are passed through the Q-Former (Dai et al., 2023; Li et al., 2023) to generate an instruction-attended query embeddings that contains the image information, and we concatenate them together to use them as few-shot samples.
A.4 Construction of the Instruction-Tuning Mixture
To evaluate the validity of our proposed benchmark, Finer, we construct an instruction-tuning mixture on top of LLaVA-1.5’s instruction-tuning mixture. We use the iNaturalist-2021 dataset to construct the mixture, sampling 20,000 instances from each of the fine-grained instances and their attributes into the training data. We structure each instruction-tuning instance in the AttrSeek pipeline format, with each instruction consistin of three turns: (i) Asking the model for the superordinate concept category, (ii) asking the model to generate a set of external, descriptive visual attributes of the superordinate concept, and finally (iii) predicting the fine-grained concept category given the superordinate concept and the generated attributes set. For each of the three steps, we use GPT-4V to generate 15 possible paraphrases of the instruction in order to avoid biasing the model to specific textual instructions and to retain the model’s instruction-following ability. We trained the model for 1 epoch, which consists of 12k steps; the checkpoint at 12k steps is the one we used to evaluate the fine-grained classification performance in Table 5.
Appendix B Prompts
In this section, we provide the input prompts used in each of the experiments, including the fine-grained visual classification and the elicitive prompting of LVLMs in Section 3.2, probing for concept-related parametric knowledge 4.1, attribute generation 4.2
B.1 Fine-Grained Visual Classification Prompts
We structure our prompts as shown in Table 8. For datasets with less than 100 class categories, we provide them along with the instruction, allowing the models to choose from the provided list of classes. Therefore, for superordinate-levels and certain coarse-levels, we provide the categories as lists so that the models solve the class generation problem by choosing from the input prompt; this is analogous to a multiple choice setting. However, for fine-grained classes, it is difficult to feed in all the concept categories in the input prompt, since some of the datasets like the iNaturalist-2021 has 10,000 categories to choose from. In order to confine the generation scope for the fine-grained labels, we decided to input the coarse-level label as denoted in Table 8, to condition the generation of the fine-grained output within a specified category space.
Dataset:iNaturalist-2021 Superordinate-level
What is the name of the organism that appears in this image? Provide your answer after "Answer:" from one of the following categories: [’Arachnids’, ’Mammals’, ’Reptiles’, ’Animalia’, ’Mollusks’, ’Plants’, ’Amphibians’, ’Ray-finned Fishes’, ’Birds’, ’Insects’, ’Fungi’].
Coarse-level
What is the name of the {concept_placeholder} that appears in this image? For example, if it’s a picture of a bengal tiger, give a coarse-grained label for the image ’Tiger’. Provide your answer after "Answer:".
Fine-level
What is the name of the {concept_placeholder} that appears in this image? For example, if it’s a picture of a bengal tiger, give a fine-grained label for the image ’Bengal Tiger’ or use its binomial nomenclature ’Panthera tigris tigris’. Provide your answer after "Answer:".
Dataset:FGVC-Aircraft Superordinate-level
What is the name of the object that appears in this image? Provide your answer after "Answer:" from one of the following categories: [’Airplane’, ’Car’, ’Train’, ’Bicycle’, ’Cell Phone’, ’Plants’, ’Dogs’, ’Birds’, ’Trucks’].
Coarse-level
What is the manufacturer of the {concept_placeholder} that appears in this image? Provide your answer after "Answer:" from one of the following categories: [’Embraer’, ’Lockheed Corporation’, ’Douglas Aircraft Company’, ’Cirrus Aircraft’, ’Airbus’, ’Antonov’, ’de Havilland’, ’Eurofighter’, ’Cessna’, ’Tupolev’, ’Dornier’, ’Yakovlev’, ’Panavia’, ’Robin’, ’ATR’, ’Beechcraft’, ’Dassault Aviation’, ’Fairchild’, ’McDonnell Douglas’, ’Fokker’, ’Gulfstream Aerospace’, ’Boeing’, ’Saab’, ’Canadair’, ’Lockheed Martin’, ’Supermarine’, ’Ilyushin’, ’British Aerospace’, ’Piper’, ’Bombardier Aerospace’].
Fine-level
What is the name of the airplane model made by {concept_placeholder} that appears in this image? For example, if it’s a picture of a Boeing 787 Dreamliner, give a fine-grained label for the image ’Boeing 787 Dreamliner’. Provide your answer after "Answer:".
Dataset:Stanford Dogs Superordinate-level
What is the name of the organism that appears in this image? Provide your answer after "Answer:" from one of the following categories: [’Arachnids’, ’Dogs’, ’Reptiles’, ’Mollusks’, ’Plants’, ’Amphibians’, ’Ray-finned Fishes’, ’Birds’, ’Insects’, ’Fungi’].
Coarse-level
What is the name of the {concept_placeholder} that appears in this image? For example, if it’s a picture of a Golden Retriever, give a coarse-grained label for the image ’Retriever’. Provide your answer after "Answer:".
Fine-level
What is the name of the {concept_placeholder} that appears in this image? For example, if it’s a picture of a Golden Retriever, give a coarse-grained label for the image ’Golden Retriever’. Provide your answer after "Answer:".
Table 8: Prompts for fine-grained image classification. The {concept_placeholder} is replaced with upper-level concept labels as illustrated in Figure 3.
Dataset:NABirds Superordinate-level
What is the name of the organism that appears in this image? Provide your answer after "Answer:" from one of the following categories: [’Arachnids’, ’Mammals’, ’Reptiles’, ’Animalia’, ’Mollusks’, ’Plants’, ’Amphibians’, ’Ray-finned Fishes’, ’Birds’, ’Insects’, ’Fungi’].
Coarse-level
What is the name of the {concept_placeholder} that appears in this image? For example, if it’s a picture of a Owl Parrot, give a coarse-grained label for the image ’Parrot’. Provide your answer after "Answer:".
Fine-level
What is the name of the {concept_placeholder} that appears in this image? For example, if it’s a picture of a Owl Parrot, give a coarse-grained label for the image ’Owl Parrot’. Provide your answer after "Answer:".
Dataset:CUB-200-2011 Superordinate-level
What is the name of the organism that appears in this image? Provide your answer after "Answer:" from one of the following categories: [’Arachnids’, ’Mammals’, ’Reptiles’, ’Animalia’, ’Mollusks’, ’Plants’, ’Amphibians’, ’Ray-finned Fishes’, ’Birds’, ’Insects’, ’Fungi’].
Coarse-level
What is the name of the {concept_placeholder} that appears in this image? For example, if it’s a picture of a Owl Parrot, give a coarse-grained label for the image ’Parrot’. Provide your answer after "Answer:".
Fine-level
What is the name of the {concept_placeholder} that appears in this image? For example, if it’s a picture of a Owl Parrot, give a coarse-grained label for the image ’Owl Parrot’. Provide your answer after "Answer:".
Dataset:Stanford Cars Superordinate-level
What is the name of the object that appears in this image?  Provide your answer after "Answer:" from one of the following categories: [’Airplane’, ’Car’, ’Train’, ’Bicycle’, ’Cell Phone’, ’Plants’, ’Dogs’, ’Birds’, ’Trucks’].
Coarse-level
What is the name of the {concept_placeholder} that appears in this image? Provide your answer after "Answer:" from one of the following categories: [’Sedan’, ’SUV’, ’Coupe’, ’Convertible’, ’Pickup’, ’Hatchback’, ’Van’]
Fine-level
What is the name of the {concept_placeholder} that appears in this image? For example, if it’s a picture of a 2006 Honda Civic LX Coupe, give a fine-grained label for the image ’2006 Honda Civic LX Coupe’. Provide your answer after "Answer:".
Table 8: Prompts for fine-grained image classification. The {concept_placeholder} is replaced with upper-level concept labels as illustrated in Figure 3.
B.2 Knowledge Probing Prompts
The knowledge probing prompts are shown in Table 9. We structure the prompt as explained in Section 4.1, where we input the Web-extracted textual attributes along with the coarse-level label for the text-only setting. Since LVLMs are good at identifying the superordinate and coarse-level concepts, we also provide the coarse-level labels as a prior for the text-only setting for a fair comparison in the analysis for fine-grained concept knowledge in the parametric knowledge space.s
B.3 Attribute Generation Prompts
The attribute generation prompts are shown in Table 10. We divide the attribute generation prompts to three different types: (i) Prompt that generates the Web-extracted attributes given the Wikipedia API retrieved concept documents, (ii) prompt that generates the attributes straight from the model given a text-only input, (iii) prompt that generates the attributes from the model given an image-only input. Note that the prompt variance between the text-only input and image-only input is intentionally minimized, i.e., minimal change in the input prompts, to more accurately isolate the effect of change in the input modalities. The prompts shown in Table 10 are used for the measuring of the modality gap experiments in Section 4.2.
B.4 Coarse-Grained Label Generation Prompts
The coarse-grained label generation prompts are shown in Table 11. We only generate the coarse-level labels for the following three datasets: (i) Stanford Dogs, (ii) Stanford Cars, (iii) CUB-200-2011, (iv) iNaturalist-2021 because they do not provide concept hierarchy like the rest of the other three datasets. For iNaturalist-2021, although the benchmark does provide the granularity hierarchy, it does so in a taxonomic manner, e.g., order, family, genus, species, which makes it challenging for the model to classify the coarse-grained categories; therefore, we generate coarse-grained labels for the dataset as well. By randomly selecting few-shot examples to guide the coarse-grained label generation, we ensure that the generative model, in this case GPT-4V, sticks to the generation of a coarse-grained label. For the faithfulness of the generated coarse-grained labels, we manually evaluate them for datasets other than iNaturalist-2021. For iNaturalist-2021, there are 1,103 coarse-grained categories, which makes it challenging to evaluate each generated coarse-grained label. We therefore group them together with their corresponding family-level category, which serves as a grouping category for the coarse-grained labels. For instance, for Euphaea fraseri, we place its coarse-grained labels, Damselfly and Dragonfly under Euphaeidae. By doing so, we not only provide room for more encompassing coarse-grained prediction, e.g., Damselflies are also classified as Dragonflies, but also distinguish the granularity setting from the fine-grained level, which requires a more specific categorization of a given species.
Dataset:iNaturalist-2021 Knowledge Probe Prompt
Can you guess the specific name (specific epithet) of an organism in the following taxonomic category given its physical attributes? Provide your answer after "Specific Epithet:".
Physical Attributes: {attribute_placeholder}
Supercategory: {supercategory_placeholder}
Kingdom: {kingdom_placeholder}
Phylum: {phylum_placeholder}
Class: {class_placeholder}
Order: {order_placeholder}
Family: {family_placeholder}
Genus: {genus_placeholder}
Specific Epithet:
Dataset:FGVC-Aircraft Knowledge Probe Prompt
Can you guess the specific name (specific type) of an Airplane in the following taxonomic category given its physical attributes? Provide your answer after "Specific Airplane:".
Physical Attributes: {attribute_placeholder}
Supercategory: {supercategory_placeholder}
Coarse-grained Category: {coarse_placeholder}
Specific Airplane:
Dataset:Stanford Dogs Knowledge Probe Prompt
Can you guess the specific name (specific type) of a Dog in the following taxonomic category given its physical attributes? Provide your answer after "Specific Dog:".
Physical Attributes: {attribute_placeholder}
Supercategory: {supercategory_placeholder}
Coarse-grained Category: {coarse_placeholder}
Specific Dog:
Table 9: Prompts for concept-related knowledge probing. The {concept_placeholder} is replaced with upper-level concept labels as illustrated in Figure 3.
Attribute Gen. (Text-Only)
What are useful visual features for distinguishing {concept_placeholder} in a photo? Provide the answer as lists of required and likely attributes. For example, for a bengal tiger (Felis Tigris) you might say:
Required:
- yellow to light orange coat
- dark brown to black stripes
- black rings on the tail
- inner legs and belly are white
- 21 to 29 stripes
Likely:
- lives in mangrove, wooded habitat
- amber, yellow eyes
- large, padded paws
- long tail
- stout teeth
’Required’ attributes are a set of external, physical attributes that allows a human to distinguish it from other similar looking concepts.
’Likely’ attributes are a set of attributes that may or may not be visible or are not one of the most discriminative features of the concept.
In the required (Required:) set, do not include relative, non-visual attributes like size or weight, only the external, visually distinguishable attributes.
Provide your response in the above format, saying nothing else. If there are no useful visual features, simply write "none".
Attribute Gen. (Wikipedia Doc; Web-Extracted)
What are useful visual, external features for distinguishing {concept_placeholder} in a photo? Given an input document (Document:) that may talk about {concept_placeholder}, provide the answer as lists of required and likely attributes. For example, for a bengal tiger (Felis Tigris) you might say:
Required:
- yellow to light orange coat
- dark brown to black stripes
- black rings on the tail
- inner legs and belly are white
- 21 to 29 stripes
Likely:
- lives in mangrove, wooded habitat
- amber, yellow eyes
- large, padded paws
- long tail
- stout teeth
’Required’ attributes are a set of external, physical attributes that allows a human to distinguish it from other similar looking concepts.
’Likely’ attributes are a set of attributes that may or may not be visible or are not one of the most discriminative features of the concept.
In the required (Required:) set, do not include relative, non-visual attributes like size or weight, only the external, visually distinguishable attributes.
If no document is given, generate from what you already know about {concept_placeholder}.
Provide your response in the above format, saying nothing else. If there are no useful visual features, simply write "none".
Table 10: Prompts for Attribute Generation. The {concept_placeholder} is replaced with coarse-grained concept labels for Image-only case and Web-Extracted cases; for Text-only case, use the fine-grained concept label since we want to extract the attributes stored in the parametric knowledge by using the fine-grained concept label as a query.
Attribute Gen. (Image-Only)
What are useful visual features for distinguishing the {concept_placeholder} in the photo? Provide the answer as lists of required and likely attributes. For example, for a bengal tiger (Felis Tigris) you might say:
Required:
- yellow to light orange coat
- dark brown to black stripes
- black rings on the tail
- inner legs and belly are white
- 21 to 29 stripes
Likely:
- lives in mangrove, wooded habitat
- amber, yellow eyes
- large, padded paws
- long tail
- stout teeth
’Required’ attributes are a set of external, physical attributes that allows a human to distinguish it from other similar looking concepts.
’Likely’ attributes are a set of attributes that may or may not be visible or are not one of the most discriminative features of the concept.
In the required (Required:) set, do not include relative, non-visual attributes like size or weight, only the external, visually distinguishable attributes.
Provide your response in the above format, saying nothing else. If there are no useful visual features, simply write "none".
Table 10: Prompts for Attribute Generation. The {concept_placeholder} is replaced with coarse-grained concept labels for Image-only case and Web-Extracted cases; for Text-only case, use the fine-grained concept label since we want to extract the attributes stored in the parametric knowledge by using the fine-grained concept label as a query.
Dataset:Stanford Cars
Generate a coarse-grained label for the following fine-grained car types.
The coarse-grained car types are as follows: ["sedan", "SUV", "coupe", "convertible", "pickup", "hatchback", "van"].
For example, if the car is a "Ford F-150 Regular Cab 2012" generate "pickup", and if the car is "Chrysler 300 SRT-8 2010", generate "sedan".
Output format is as follows:
Fine-grained Car Name: Ford F-150 Regular Cab 2012
Car Name: pickup
Fine-grained Car Name: Chrysler 300 SRT-8 2010
Car Name: sedan
Fine-grained Car Name: Hyundai Santa Fe 2008
Car Name: SUV
Dataset:CUB-200-2011
Generate a coarse-grained label for the following fine-grained bird types.
For example, if the bird is a "bald eagle (Haliaeetus leucocephalus)" generate "Eagle", and if the bird is "Pine grosbeak", generate "Finch".
Output format is as follows:
Fine-grained Bird Name: Bald eagle
Bird Name: Eagle
Fine-grained Bird Name: Pine grosbeak
Bird Name: Finch
Fine-grained Bird Name: The black backed woodpecker
Bird Name: Woodpecker
Dataset:Stanford Dogs
Generate a coarse-grained label for the following fine-grained dog types.
For example, if the dog is a "Cavalier King Charles Spaniel" generate "Spaniel", and if the dog is "The Dear-Headed Chihuahua", generate "Chihuahua".
Output format is as follows:
Fine-grained Dog Name: Cavalier King Charles Spaniel
Dog Name: Spaniel
Fine-grained Dog Name: Curly-coated retriever
Dog Name: Retriever
Fine-grained Dog Name: Newfoundland
Dog Name: Newfoundland
Table 11: Prompts for Coarse-grained Label Generation. We generate the coarse-grained labels for each of the dataset, in case the dataset does not provide the concept hierarchy. The few-shot sampled are randomly sampled from the training instances.